OpenStudy (anonymous):

A store offers a 15% discount on a skirt. If Bernice pays $40.80 for the skirt, what was the list price of the skirt before the discount? $61.00 $48.00 $44.61 $40.10 $31.00 first off who would pay 40 something dollars for a skirt an second i cant figure this out ive tried 3 times and it comes out with $46.92 can someone help me and show me wht im doing wrong?

OpenStudy (igreen):

We can make an equation. \(\sf x - (x \cdot 15\%) = 40.80\) Can you solve that for 'x'?

OpenStudy (anonymous):

ok and yea i could

OpenStudy (igreen):

Alright, tell me what you come up with.

OpenStudy (anonymous):

ummmm whered the subtractng x come from?

OpenStudy (igreen):

When taking 15% off, we multiply 15% to the original price(x) and then subtract that to the original price(x). So our equation is multiply 15% to 'x', and subtracting that from 'x', and it equals $40.80.

OpenStudy (anonymous):

ok tht makes more since than wht u just put as the equation

OpenStudy (igreen):

Can you simplify \(\sf x \cdot 15\%\)?

OpenStudy (anonymous):

yea

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

61*15% is 9.15 so i subtrct tht from 61 or 40.80?

OpenStudy (igreen):

Where did you get 61 from?

OpenStudy (anonymous):

the answers

OpenStudy (anonymous):

lol

OpenStudy (igreen):

We want to simplify \(\sf x \cdot 15\%\) \(\sf 15\% = 0.15\) \(\sf x \cdot 0.15 = 0.15x\) Understand?

OpenStudy (jaceykilla):

igreen help me w my question pleaseeee

OpenStudy (anonymous):

well ya i knew how to multiply with percents i just dotn get the question

OpenStudy (anonymous):

im leaving in like 2 minutes so :/

OpenStudy (igreen):

That gives us: \(\sf x - 0.15x = 40.80\) 1 - 0.15 = ?

OpenStudy (anonymous):

1-0.15=0.85

OpenStudy (igreen):

Right, so we have: \(\sf 0.85x = 40.80\) Divide 0.85 to both sides

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

x=48

OpenStudy (anonymous):

OMG thank u!

OpenStudy (igreen):

Yes, that's correct.

OpenStudy (anonymous):

i gtg